# MindGYM: What Matters in Question Synthesis for Thinking-Centric Fine-Tuning?

### v2

@@ Abstract
11footnotetext: Equal contribution.
@@ 1 Introduction
We validate MindGYM on six reasoning-intensive benchmarks across multiple vision-language models (VLMs) architectures. As shown in Table 1, with only 400 synthetic samples, our method yields significant gains: for example, our method boosts performance of Qwen2.5-VL-7B qwen25vl by over 16% on MathVision-Mini. These improvements are consistent across models of varying scales and architectures, demonstrating the generality of our approach. Further analysis via Data-Juicer dj2 reveals that on Qwen2.5-VL-32B qwen25vl , our synthetic data achieves a +16.7% average quality improvement over baselines and a 67.91% reduction in quality variance. This highlights an important insight: not only does cognitively guided synthesis yield higher-quality data, but stability in quality—as measured by low variance—is critical for consistent fine-tuning. Moreover, beyond the high-quality textual data synthesis presented in the main page, we also explore MindGYM’s scalability to multimodal settings in Appendix C, demonstrating its effectiveness and potential to both VLMs and large language models (LLMs).
@@ 2 Related Works
Synthetic data generation has long been a tool for improving model reasoning. In language-only domains, question synthesis approaches have emphasized structure, difficulty scaling, and multi-step deduction, particularly in math or logic domains lewkowycz2022solving ; shao2024deepseekmath . However, most existing methods either depend on rigid templates or domain-specific structures, which constrains their generalizability. Multimodal question synthesis introduces further complexity due to the need for cross-modal reasoning and abstract concept grounding. Methods like Visual Program Induction pmlr-v162-duan22c or GRILL parcalabescu2022valse attempt structured generation but often produce shallow or repetitive questions. Moreover, many pipelines prioritize syntactic fluency over cognitive depth, failing to elicit the higher-order thinking needed for robust generalization. This motivates the need for synthesis frameworks that directly embed cognitive traits—such as reasoning breadth, depth, and progression—into the data itself.
@@ 3 The Proposed MindGYM
Figure 1: The proposed MindGYM framework incorporates a cognitively guided data synthesis pipeline with four stages: ① Context Generation, ② Single-Hop Question Synthesis, ③ Multi-Hop Composition with Thinking Trace, and ④ Structured Extraction. Starting from a meta-topic and guided by a shared cognitive objective, the model iteratively builds background context, atomic reasoning steps, composite questions with interpretable thinking traces, and final structured QA samples for downstream use. Color-coded regions and arrows illustrate the hierarchical progression from simple reasoning to advanced problem-solving and emphasizing conceptual depth.
@@ 3.2 Seed Single-Hop Question Synthesis
To ensure reasoning diversity and depth, we construct eight orthogonal meta-topics derived from established cognitive theories, including unified cognitive architectures newell1994unified and dual-representation models sun2001duality . These meta-topics span four core reasoning dimensions: quantitative, causal, temporal, and social-ethical, offering a comprehensive blueprint for cognitively rich question generation. Detailed meta-topics are provided in Appendix B.2.
@@
To enhance diversity in seed question generation, MindGYM further applies semantic vectorization to textual components of data in the synthesis pool. We use an empirically determined cosine similarity threshold to evaluate the overlap between newly generated and existing samples. If similarity exceeds the threshold, the system initiates a regeneration cycle to eliminate semantically redundant content, continuing until the seed question pool reaches the target size $N$. This reject sampling mechanism widens the breadth of reasoning in synthetic data.
@@ 3.3 Challenging Multi-Hop QA Synthesis
Given the seed questions $\{q^{j}_{i}\}_{j=1}^{k}$ from Section 3.2, we treat them as updated context $P_{\text{context}}^{\text{single}}$ for this phase. The model $\pi$ is now prompted with the triplet $(P_{\text{cog}},P_{\text{context}}^{\text{single}},P_{\text{task}}^{\text{multi-hop}})$ to synthesize a more complex, cognitively grounded question $Q_{i}$. The task instruction $P_{\text{task}}^{\text{multi-hop}}$ specifies an abstract reasoning requirement (e.g., comparing entities, bridging concepts, or analyzing temporal relations), which guides the model to self-compose from prior seeds. Details of composition types are provided in Appendix B.
@@ 4 Experiments
To comprehensively evaluate the performance of our proposed synthetic data method, we conduct extensive experimental validation on multiple representative evaluation sets. Full descriptions of datasets, benchmarks, implementation details, and baselines are provided in Appendix A.
@@ 4.2 Main Results
Despite using only 400 synthetic training samples, MindGYM achieves performance that not only matches but exceeds baselines trained with 10$\times$ more data. This includes outperforming Open-O1 (4k) in both textual and multimodal tasks. On MathVision, MindGYM reaches 28.62 while Open-O1 (4k) collapses to 17.43. This stark contrast underscores the high quality and utility of the data synthesized by MindGYM.
@@ 4.3 Data analysis
We establish a two-dimensional evaluation framework leveraging GPT-4 as an expert scorer. The protocol assesses: (1) thinking Depth via derivation complexity, and (2) thinking Breadth by solution diversity. Each metric operates under a standardized scoring system, with detailed specifications provided in Appendix B.2.5. Comparative evaluations against four baselines are conducted on the MathVision dataset as a benchmark. The experimental results documented in Table 3 demonstrate MindGYM’s consistent superiority across both evaluation axes, particularly achieving 15.8% average improvements in thinking depth over the raw model and other competitive baselines.
@@ 4.4 Ablation studies
Ablation of Stream of Consciousness.
Table 4: Ablation study results of removing stream of consciousness (w/o SC), utilizing English in data synthesis (Syn-EN), changing the order of our fine-tuning (w/o OF) steps and Relation Balanced (with RB), complete results detailed in Appendix D.
Datasets MM-Avg Text-Avg Avg
@@
We conduct the ablation study by removing the stream of consciousness module from our prompt while maintaining the same fine-tuning paradigm for training and evaluation. The experimental results, as shown in Table 4, indicate that the removal of the stream of consciousness module leads to a decrease across all Avg scores, with reduction of 1.48 points. This decline highlights the significance of the stream of consciousness module in our approach, underscoring its essential role in enhancing the depth of thinking and overall performance of the model.
@@
Linguistic Difference.
Based on the selected Qwen qwen25vl , we utilize Chinese as the language for data synthesis. Additionally, we explore the effects of using English for synthesis, as shown in Table 4 under the Syn-EN condition. The results indicate the performance degradation when using English, with a decrease of 1.31 points. Beyond the inherent characteristics of Qwen, we hypothesize that the higher information density of Chinese makes it beneficial for the model’s thinking learning.

### v3

@@ Abstract
33footnotetext: FSIETP: Guangdong Provincial Key Laboratory of Fire Science and Intelligent Emergency Technology.
@@ 1 Introduction
We validate MindGYM on six reasoning-intensive benchmarks across multiple vision-language models (VLMs) architectures. As shown in Table 1, with only 400 synthetic samples, our method yields significant gains: for example, our method boosts performance of Qwen2.5-VL-7B qwen25vl by over 16% on MathVision-Mini. These improvements are consistent across models of varying scales and architectures, demonstrating the generality of our approach. Further analysis via Data-Juicer dj2 reveals that on Qwen2.5-VL-32B qwen25vl , our synthetic data achieves a +16.7% average quality improvement over baselines and a 67.91% reduction in quality variance. This highlights an important insight: not only does cognitively guided synthesis yield higher-quality data, but stability in quality—as measured by low variance—is critical for consistent fine-tuning. Compared with CoT wei2022chain and ToT yao2023tree baselines, our multi-step cognitive reasoning framework demonstrates superiority. We further observe that Chinese-synthesized data outperform English and mixed-language variants. Moreover, beyond the high-quality textual data synthesis presented in the main page, we also explore MindGYM’s scalability to multimodal settings in Appendix C, demonstrating its effectiveness and potential to both VLMs and large language models (LLMs).
@@ 2 Related Works
Synthetic data generation has long been a tool for improving model reasoning. In language-only domains, question synthesis approaches have emphasized structure, difficulty scaling, and multi-step deduction, particularly in math kuang2025atomic or logic domains lewkowycz2022solving ; shao2024deepseekmath . However, most existing methods either depend on rigid templates or domain-specific structures, which constrains their generalizability. Multimodal question synthesis introduces further complexity due to the need for cross-modal reasoning and abstract concept grounding. Methods like Visual Program Induction pmlr-v162-duan22c or GRILL parcalabescu2022valse attempt structured generation but often produce shallow or repetitive questions. Moreover, many pipelines prioritize syntactic fluency over cognitive depth, failing to elicit the higher-order thinking needed for robust generalization. This motivates the need for synthesis frameworks that directly embed cognitive traits—such as reasoning breadth, depth, and progression—into the data itself.
@@ 3 The Proposed MindGYM
Figure 1: The proposed MindGYM framework incorporates a cognitively guided data synthesis pipeline with four stages: Context Generation, Single-Hop Question Synthesis, Multi-Hop Composition with Thinking Trace, and Structured Extraction. Starting from a meta-topic and guided by a shared cognitive objective, the model iteratively builds background context, atomic reasoning steps, composite questions with interpretable thinking traces, and final structured QA samples for downstream use. Color-coded regions and arrows illustrate the hierarchical progression from simple reasoning to advanced problem-solving and emphasizing conceptual depth.
@@ 3.2 Seed Single-Hop Question Synthesis
To ensure reasoning diversity and depth, we construct eight orthogonal meta-topics derived from established cognitive theories, including unified cognitive architectures newell1994unified and dual-representation models sun2001duality . These meta-topics span four core reasoning dimensions: quantitative, causal, temporal, and social-ethical, offering a comprehensive blueprint for cognitively rich question generation. Detailed meta-topics are provided in Appendix A.2.
@@
To enhance diversity in seed question generation, MindGYM further applies semantic vectorization to textual components of data in the synthesis pool ling2025diversity . We use an empirically determined cosine similarity threshold to evaluate the overlap between newly generated and existing samples. If similarity exceeds the threshold, the system initiates a regeneration cycle to eliminate semantically redundant content, continuing until the seed question pool reaches the target size $N$. This reject sampling mechanism widens the breadth of reasoning in synthetic data.
@@ 3.3 Challenging Multi-Hop QA Synthesis
Given the seed questions $\{q^{j}_{i}\}_{j=1}^{k}$ from Section 3.2, we treat them as updated context $P_{\text{context}}^{\text{single}}$ for this phase. The model $\pi$ is now prompted with the triplet $(P_{\text{cog}},P_{\text{context}}^{\text{single}},P_{\text{task}}^{\text{multi-hop}})$ to synthesize a more complex, cognitively grounded question $Q_{i}$. The task instruction $P_{\text{task}}^{\text{multi-hop}}$ specifies an abstract reasoning requirement (e.g., comparing entities, bridging concepts, or analyzing temporal relations), which guides the model to self-compose from prior seeds. Details of composition types are provided in Appendix A.
@@ 4 Experiments
To comprehensively evaluate the performance of our proposed synthetic data method, we conduct extensive experimental validation on multiple representative evaluation sets. Full descriptions of datasets, benchmarks, implementation details, and baselines are provided in Appendix B.
@@ 4.2 Main Results
[…]
@@ 4.3 Data analysis
We establish a two-dimensional evaluation framework leveraging GPT-4 as an expert scorer. The protocol assesses: (1) thinking Depth via derivation complexity, and (2) thinking Breadth by solution diversity. Each metric operates under a standardized scoring system, with detailed specifications provided in Appendix A.2.5. Comparative evaluations against four baselines are conducted on the MathVision dataset as a benchmark. The experimental results documented in Table 3 demonstrate MindGYM’s consistent superiority across both evaluation axes, particularly achieving 15.8% average improvements in thinking depth over the raw model and other competitive baselines.
@@ 4.4 Ablation studies
Impact of Structured Cognitive Thinking.
Table 4: Ablation study results of removing structured cognitive thinking (w/o SC), utilizing English in data synthesis (Syn-EN), changing the order of our fine-tuning (w/o OF) steps and Relation Balanced (with RB), structured cognitive thinking with CoT and ToT synthes (CoT and ToT), complete results detailed in Appendix D.
Datasets $M$-Avg Text-Avg Avg
@@
CoT 59.74 51.98 55.86 $\downarrow$
ToT 59.78 52.07 55.92 $\downarrow$
We investigate the role of the structured cognitive thinking (SC) module, which models multi-step reasoning behaviors such as planning, divergent exploration, and selective consolidation. This module is implemented through our multi-step synthesis scheme, where question and answer generation are explicitly decomposed into reasoning stages. To contrast with this design, we evaluate two widely used reasoning paradigms—Chain-of-Thought (CoT) and Tree-of-Thought (ToT)—which directly generate multi-hop questions without intermediate reasoning supervision. As shown in Table 4, MindGYM-Text achieves the highest average score of 57.84, outperforming both CoT (55.86) and ToT (55.92). This demonstrates that multi-step structured synthesis provides richer reasoning signals than direct single-pass generation. Moreover, removing SC (MindGYM-Text w/o SC) results in a performance drop to 56.36, confirming that structured cognitive guidance itself contributes to the improvement. Overall, these results suggest that multi-step, cognition-guided synthesis is more effective than direct CoT or ToT prompting for constructing high-quality reasoning data.
Table 5: GPT-4o-based qualitative evaluation of synthesized data.
logically flawed syntactically ambiguous correctness hallucination
false 328 394 357 372
true 71 5 42 27
Effect of Data Quality and Filtering.
We further analyze the quality and impact of our data through a controlled qualitative and filtering study. The dataset is evaluated by GPT-4o along four dimensions: logical consistency, linguistic clarity, factual correctness, and hallucination tendency. As shown in Table 5, over 89% of samples are judged factually correct, 93% non-hallucinatory, and fewer than 1% syntactically ambiguous, demonstrating the overall linguistic and logical soundness of the MindGYM synthesis process.
Table 6: Impact of filtering on model performance.
Model text-avg $M$-avg Overall Avg
MindGYM (full) 61.27 54.42 57.84
MindGYM (clean) 61.08 53.89 57.49
To assess whether further cleaning benefits model training, we remove samples flagged as logically flawed or hallucinatory by GPT-4o and retrain the model on this “cleaned” subset. As presented in Table 6, the cleaned data lead to slightly lower performance compared to the full dataset (57.49 vs. 57.84). This finding indicates that the effectiveness of MindGYM does not solely stem from factual accuracy, but rather from the rich reasoning trajectories embedded in the synthesized data. Even imperfect samples often encode meaningful intermediate reasoning steps—such as decomposition, hypothesis formation, and evidence evaluation—that contribute to more generalizable reasoning behavior. These results suggest that, for thinking-centric instruction tuning, maintaining reasoning diversity and cognitive structure is more beneficial than enforcing strict factual precision.
Effect of Multimodal Synthesis and Vision Backbone Freezing.
Table 7: Ablation on multimodal synthesis and vision backbone freezing. “CN” = Chinese text-only synthesis; “Flux” = additional text-to-image generation using Flux; “Frozen/Unfrozen” indicates whether the vision encoder is trainable.
Setting text-avg $M$-avg Avg
CN (text only, Frozen) 61.27 54.42 57.84
Flux (image gen., Frozen) 60.91 52.68 56.80
CN (text only, Unfrozen) 60.40 52.30 56.35
We assess the influence of multimodal synthesis and vision backbone freezing. As shown in Table 7, the default Chinese text-only setting with a frozen encoder achieves the best overall average of 57.84. Introducing Flux-based image generation slightly reduces performance to 56.80, while unfreezing the vision backbone leads to a further drop to 56.35. These results indicate that additional multimodal synthesis and vision unfreezing introduce minor instability without clear benefit, validating our design choice of a text-dominant and frozen-encoder configuration for stable reasoning performance. For more information about Flux dataset synthesis, see the Appendix A.2.6.
Linguistic Difference.
Table 8: Cross-lingual comparison of MindGYM data synthesis.
Setting text-avg $M$-avg Avg
CN (Chinese only) 61.27 54.42 57.84
MIX (CN + EN) 61.16 53.51 57.34
EN (English only) 60.30 52.77 56.50
Based on the selected Qwen qwen25vl , we examine the impact of synthesis language using three variants of MindGYM data: Chinese (CN), mixed Chinese–English (MIX; 50% Chinese and 50% English), and English (EN), as shown in Table 8. The Chinese-only configuration achieves the highest overall performance with an average score of 57.84 and is adopted as our default setting. In contrast, using English for synthesis leads to a noticeable performance drop of 1.34 points on average, while the mixed-language variant (MIX) also performs slightly worse than CN. These findings indicate that Chinese synthesis yields more effective reasoning supervision, likely benefiting from its higher information density and compact semantic structure, which align better with the thinking-centric nature of our data construction.
@@ 5 Conclusion & Future Works
Acknowledgement
This research is supported by Key-Area Research and Development Program of Guangdong Province, Granted No. 2024B1111060004.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction accurately summarize the paper’s main contributions and scope, which are consistently developed and supported throughout the paper (see Sec. 3 and Sec. 4.2).
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Section 3.4 discusses the limitations of our approach and outlines potential directions for future work to extend our findings.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: This work focuses on developing a data synthesis method to generate high-quality training data, with primary emphasis on practical implementation rather than theoretical analysis.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We propose the details of our data synthesis method in Sec. 3, with reproducibility details in Sec. 4.1 and Appendix B.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: All code and data are released to shed light on further research and applications.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We present the details of experimental setting in Sec. 4.1 and Appendix B.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: Due to the considerable costs involved in training hundreds large language models, we run each experiment once and conduct extensive experiments across multiple benchmarks using different models to observe detailed results, which are reported in Sec. 4.2 and  4.3.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Experimental compute resources are provided in Appendix B.2.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: This research fully complies with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Our societal impact and broader implications are discussed in Sec. 3.4
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Our study involves neither high-risk data nor models.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All models and datasets used in our experiments (Sec. 3 & Sec. 4.1) are publicly available, with proper attribution and adherence to their respective licenses.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: All code and data are open-sourced. See Section 3 for synthesis methods and Section 3.5 for usage instructions.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: Our data synthesis method employs self-supervised model generation, requiring no human intervention.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: No human subjects were involved in this study. All data were synthetically generated through self-supervised model training without human participation.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.